# Quantum-centric Supercomputing for Materials Science: A Perspective on Challenges and Future Directions

### v1

@@ II.4 Problem-level preprocessing and optimization
Despite these recent advances in computational methods for modeling electronic structures, significant research challenges exist, particularly involving the treatment of “larger” active spaces and the inherent poor scaling of multireference perturbation theories. Numerous research efforts have been directed towards the treatment of larger active spaces using classical electronic structure techniques like DMRG, selected CI, and quantum Monte Carlo methods. However, expanding the active space relies on physically motivated approximations to balance computational cost [7, 235, 236, 237, 9]. The basic philosophy in all these approaches is focusing on the most important configurations while eliminating less significant ones, based on physically motivated criteria. Quantum computing holds the promise of effectively modeling these larger active spaces accurately, especially as the number of available qubits increases and noise correction techniques become more accurate. Embedding techniques such as LAS-UCC may help accelerate this process. While perturbation theories scale poorly on classical hardware, recent studies on formulating quantum implementations of perturbation theories[238] highlight the potential for creating multireference quantum algorithms, including NEVPT2-DMET. With more qubits being available, the primary research challenge still remains to determine the optimal methods for quantum implementation that maintain a favorable balance between computational cost and the desired accuracy.
@@
Our second example of problem-aware preprocessing is VQE ansatz design [33]. VQE (and its variants) is the most widely used algorithm on quantum computers for the eigenvalue problem, although its experimental applications to date have been limited to classically-simulable demonstrations (e.g., [31, 32, 33, 34, 35, 36, 37, 38, 39, 40, 41, 42, 43, 44, 45, 46, 47, 49, 51, 52]). The core of the VQE algorithm is the parameterized quantum circuit $U(\vec{\theta})$, usually called the ansatz, whose output is a manifold of states that should contain an approximation to the ground state of the system. The ansatz can be evaluated according to two key attributes: expressivity and practicality. The former measures how well the ansatz can approximate the solution to the problem to a certain accuracy, while the latter takes circuit depth, the difficulty of optimizing those parameters, and the qubit connectivity of the underlying hardware into consideration. Quantum advantage in VQE requires identifying an ansatz that satisfies both criteria, i.e., it is a feasible circuit on a given device, it is trainable at scale on a classically hard problem, and it can represent a low-energy state with sufficient accuracy to surpass classical methods.
@@
Two extreme examples of ansatze are the unitary coupled-cluster (UCC) ansatz [33, 36, 48, 239] and hardware efficient ansatzes [38]. The UCC is chemically-inspired and thus very expressive for most chemistry and materials but usually entails a large number of gates and is likely beyond near-term hardware. The hardware-efficient ansatzes are problem-agnostic and designed specifically to use the native gates and connectivity of a given device, but may be impractical to optimize due to the barren plateau problem [240], and also may fail to be sufficiently expressive at low depth, lead to symmetry breaking, and non-smooth potential energy surfaces [241].
@@
Efforts have been put into designing new ansatzes to strike a balance between these two dimensions and unify the advantages of these two types of ansatz [242, 243, 244, 245]. The qubit coupled-cluster ansatz [242, 246, 247] and ADAPT-VQE [245] are among such examples, where a pre-screening process is proposed to select certain Pauli operators to generate the gates in the ansatz, thereby reducing the circuit depth and number of parameters. These methods that trim the number of variational parameters are promising, but there is still a lack of evidence for the scalability of parameter training. An exception may be HMP2 [50], where second-order perturbation theory is employed to reduce the number of variational parameters, while providing real-time, quantum computation-based methodologies to identify additional Pauli operators to include in the VQE progression. These advancements indeed merit further research despite outstanding challenges, because of the simplicity of variational quantum algorithms in all other respects (particularly in circuit complexity) and the extensive success of classical variational algorithms.
@@
Another challenge with ADAPT-VQE-like algorithms lies in the larger measurement overhead due to the selection of Pauli operators, which can however be mitigated by the use of IC-POVM measurements [248]. While it is important to design an ansatz structure that is most appropriate for a target problem, appropriately initializing the ansatz parameters is an equally critical task when it comes to the trainability of the resulting optimization. This is because random parameter initialization can be extremely far from the optimal parameters, and the ability for VQE to converge to the optimal parameters (and thereby find the ground state) will heavily depend on the optimizer and the effects of noise. Identifying good classical initialization can therefore considerably improve the accuracy and speed of VQE convergence. One promising method is CAFQA [249]. The CAFQA ansatz is built with only Clifford gates and is therefore classically simulable in polynomial time. A discrete optimizer is used to search through different possible Clifford gate parameters until the optimizer converges to a set of parameters that minimizes the VQE expectation (to the best possible extent within the Clifford space). This ‘CAFQA’ state is then used to initialize traditional VQE. There is an opportunity to explore other forms of classical simulation to bootstrap VQE—near-Clifford circuits and tensor network approaches are promising directions [250, 251]. All these methods could benefit from problem-specific optimization. For instance, restricting the CAFQA Clifford search to a state space that is expected to have some overlap with the ground state can improve initialization speed and accuracy.
II.5 Classical data loading
@@
At the highest level, quantum computation involves the following steps: first, we establish a classical description for a specific problem. This description is input into a quantum circuit. After executing the quantum circuit, we retrieve classical measurement results as the solution to the problem. Within this workflow, an essential challenge is the data input. A robust quantum algorithm must explicitly detail this input technique and the associated hardware.
@@
For some types of use cases, such as many quantum simulation algorithms, the problems themselves are quantum mechanical in nature, and often this permits direct mappings to quantum circuits, such as those discussed in Sec. II.1. In more general cases, including when the inputs are classical data, scientists who focus on theoretical complexity analysis often abstract such input methods as an ‘oracle’ and investigate how many times this oracle must be applied, using this as a basis to estimate the algorithm’s complexity. However, creating a physical realization of a quantum oracle is also a fundamental aspect of actualizing quantum computing.
@@
Quantum Random Access Memory (QRAM) [252] serves as the physical embodiment of the quantum oracle. Unlike traditional RAM, where a single piece of data can be swiftly loaded into the central processor, QRAM allows for multiple classical data items to be loaded concurrently in superposition. This unique capability positions QRAM as a bridge between classical and quantum computing realms. More than that, QRAM can function as a rapid and versatile quantum oracle, effectively addressing the data-input bottleneck prevalent in numerous quantum applications. QRAM is also a fundamental assumption for many quantum algorithms. For instance, many quantum machine learning algorithms [253, 254] may require QRAM as a prerequisite to potentially achieve quantum speedups, although some algorithms have been discovered to achieve the speedups without [255]. Currently, given the significant challenges in physically constructing QRAM, both theoretical and experimental research on QRAM is an active field of study.
@@
For quantum simulation and material sciences, QRAM could primarily be used to provide fast, efficient, and reliable initial states and Hamiltonian encoding for quantum simulation of material science problems. Those QRAM architectures could provide useful interfaces, benefiting practical use cases of quantum simulation to large-scale, challenging problems in material science.
QRAM construction:
@@
  • Quantum data centers. Nowadays, data centers are important businesses in the digital society, and it is natural to have a quantum version of data centers in the coming era. Refs. [262, 263] propose a theory of quantum data centers (QDCs), where QRAM and quantum networks are combined to provide an elementary definition of QDC. These works show that QDCs with several applications in QRAM as fast quantum memories, will provide fast, secure, and accurate applications in quantum computing, communication, and sensing.
@@
Fault-tolerant designs of quantum random access memory are investigated by [267], showing that QRAM with full fault tolerance might be very challenging on a large scale. However, on the other hand, in [258], it is shown that noisy QRAM might still be error-resilient for generic setups of decoherence and noisy channels. In fact, though fanout QRAM designs might be vulnerable to decoherence, rendering them non-scalable, the bucket brigade QRAM architecture is known for its high noise resilience [252, 258]. Specifically, the infidelity of a query grows only logarithmically with memory size [258]. Yet, past studies leaned on specific noise models, raising questions about the practical advantages in real-world implementations. Ref. [258] delved into the effects of decoherence on QRAM in a comprehensive manner. A significant takeaway is the affirmation that the logarithmic infidelity scaling persists across varied error channels, including challenges like depolarizing noise and coherent errors. The research identifies limited entanglement among the memory components as the chief reason behind this noise resilience. Interestingly, this insight also suggests potential architectural simplifications without compromising noise resilience, serving an interesting future research question. This understanding implies that with contemporary hardware, QRAM might be potentially realized in real-world noisy environments, facilitating high-fidelity queries even with early fault-tolerant technologies. Additionally, when quantum error correction is integrated, the bucket brigade architecture continues to show promise, enhancing hardware efficiency and fortifying against logical errors [252, 258].
Experimental efforts towards QRAM:
@@
II.6 Circuits, transpilation, and architecture
Just as in classical computing, every quantum computation that is to be carried out on a quantum computer needs to be broken down to a netlist of native operations that admit a direct implementation on a quantum computer. One of the most widely-used representation of such a netlist is the so-called quantum circuit, which is to be transpiled for a target quantum computer backend. Different quantum hardware backend may support different qubit-to-qubit connectivity, which limits the interactions between the qubits that may be implemented natively.
To this end, quantum applications in materials science are no exceptions. Indeed, quantum circuits for Heisenberg hamiltonian simulation [151, 152], relevant for many-body localization and room-temperature superconductivity, and electronic structure calculations for small molecules [48, 50, 53], relevant for chemical reactions, have been optimized for both pre-fault tolerant and fault-tolerant settings before. Other examples include particle physics simulations [276, 130, 132], where fault-tolerant cost, including the number of non-Clifford quantum gates and qubit counts, have been minimized. In all these examples, it is critical that the resulting netlist of instructions to be carried out by a quantum computer is optimized, targeting time to solution: It is a key metric in all types of computation, including quantum. Case in point, if time to solution is of no matter, the computational advantages offered by quantum computing is moot and classical computing suffices, and a high-fidelity operation of a quantum computer would become unnecessary since repeated executions of computation can be performed until a correct answer is obtained. The important roles that circuits, transpilation, and architecture will continue to play in leveraging the computational power enabled by quantum computation then form the central topics of this section.
II.6.1 Connectivity and Architecture
We begin by highlighting key topics around qubit connectivity, a fundamental aspect of QPUs that dictates the efficiency, scalability, and feasibility of the quantum devices. We first consider SWAP networks, highlighting their role in enabling Hamiltonian simulation in quantum systems with limited connectivity. The discussion then moves to distributed quantum computing, emphasizing its importance in scaling quantum computing beyond the capabilities of current QPUs. Quantum circuit cutting is also examined for its practicality in decomposing complex quantum circuits into smaller segments, facilitating distributed processing.
@@ II.6.1 Connectivity and Architecture
One of the tools that can be leveraged to effectively implement Hamiltonian simulation primitives on hardware with minimal (linear) connectivity is the SWAP network [277, 278, 279]. Through a SWAP network (depicted in Figure 4), all $\binom{n}{2}\in O(n^{2})$ pairs of (commuting) interactions between $n$ qubits can be accomplished in just $O(n)$ circuit layers, even on a quantum computer architecture that admits nearest-neighbor qubit-to-qubit connectivity. With the exception of an all-to-all architecture that admits a single-instruction implementation of commuting, overlapping two-qubit gates [280, 281, 282], the asymptotic scaling of $O(n)$ matches that admitted by an architecture that allows for random-access two-qubit interactions that may be implemented in parallel so long as the interactions do not overlap. In the context of the Materials Science applications in this paper, any “dense” interactions with commuting terms can have a natural implementation on an all-to-all architecture (c.f. Refs. [152, 50, 132, 53]) or on a limited-connectivity architecture with SWAP networks. Reducing the overhead in moving the quantum information around for optimal implementation of a given set of two-qubit interactions remains an area of ongoing research.
@@
Just like classical HPC, distributed quantum computing is necessary for the continued scaling of quantum computing. While monolithic quantum processing units (QPUs) continue to make significant progress in both sizes and qubit qualities, there remains a huge gap between available QPUs and practical benchmarks. To make matters worse, developments from classical computing demonstrate that the computing and memory capacity gaps between large-scale industrial use cases and single computing cores only get larger, despite having more powerful monolithic cores. To combat this increasing gap, almost all industry-scale classical applications today such as Large Language Model (LLM) training, and even many personal-scale workloads such as intensive gaming, happen in parallel. In fact, distributed computing is only more imperative for quantum computing as single-core scalability poses significant engineering challenges.
There have been distributed quantum computing proposals from both hardware and software researchers. IBM, Quantinuum, and IonQ all announced plans to develop modular QPUs with interconnects to enable distributed quantum computing. Ref. [283] models hardware performances of potential superconducting modular QPUs. Ref. [284] proposes compilation algorithms for distributed QPUs with EPR pair connections. Ref. [285] proposes ClusterVQE which enables domain-specific optimizations to map applications from a target domain (here, variational quantum algorithms) to distributed QPUs. Further research across the stack from applications to devices is necessary to enable distributed quantum computing with tolerable overheads.
@@
Despite its ability to run previously intractable workloads, there is still a gap between the computational capacity of circuit cutting now and future practical quantum workloads. The first theory proof [286] proposed an exponential classical reconstruction cost. Instead, subsequent works [289] drastically reduce the overhead to sub-exponential but remain prohibitive in the worst case. Filling the gap calls for further developments in classical reconstruction techniques, algorithm-backend co-design, and quantum computing cloud backend system adaptations [290].
@@
The QRAM and QROM memory primitives discussed earlier can be optimally implemented with appropriate compiler techniques. As one example, consider the Gray Code, a sequence of the $2^{n}$ n-bit bitstrings such that only 1 bit changes between sequential neighbors, for instance, 000, 001, 011, 010, 110, 111, 101, 100. Integrating awareness of the Gray Code in a compiler can significantly reduce gate counts for memory implementations. Additional compilation techniques for these memory primitives, including routing and mapping, are discussed in prior work [291, 273, 274].

### v2

@@ III.4 Problem-level preprocessing and optimization
Despite these recent advances in computational methods for modeling electronic structures, significant research challenges exist, particularly involving the treatment of “larger” active spaces and the inherent poor scaling of multireference perturbation theories. Numerous research efforts have been directed towards the treatment of larger active spaces using classical electronic structure techniques like DMRG, selected CI, and quantum Monte Carlo methods. However, expanding the active space relies on physically motivated approximations to balance computational cost [7, 236, 237, 238, 9]. The basic philosophy in all these approaches is focusing on the most important configurations while eliminating less significant ones, based on physically motivated criteria. Quantum computing holds the promise of effectively modeling these larger active spaces accurately, especially as the number of available qubits increases and noise mitigation techniques become more powerful. Embedding techniques such as LAS-UCC may help accelerate this process. While perturbation theories scale poorly on classical hardware, recent studies on formulating quantum implementations of perturbation theories[239] highlight the potential for creating multireference quantum algorithms, including NEVPT2-DMET. With more qubits being available, the primary research challenge still remains to determine the optimal methods for quantum implementation that maintain a favorable balance between computational cost and the desired accuracy.
@@
Our second example of problem-aware preprocessing is VQE ansatz design [33]. VQE (and its variants) is the most widely used algorithm on quantum computers for the eigenvalue problem, although its experimental applications to date have been limited to classically-simulable demonstrations; see, e.g., [31, 32, 33, 34, 35, 36, 37, 38, 39, 40, 41, 42, 43, 44, 45, 46, 47, 49, 51, 52]. The core of the VQE algorithm is the parameterized quantum circuit $U(\vec{\theta})$, usually called the ansatz, whose output is a manifold of states that should contain an approximation to the ground state of the system. The ansatz can be evaluated according to two key attributes: expressivity and practicality. The former measures how well the ansatz can approximate the solution to the problem to a certain accuracy, while the latter takes circuit depth, the difficulty of optimizing those parameters, and the qubit connectivity of the underlying hardware into consideration. Quantum advantage in VQE requires identifying an ansatz that satisfies both criteria, i.e., it is a feasible circuit on a given device, it is trainable at scale on a classically hard problem, and it can represent a low-energy state with sufficient accuracy to surpass classical methods.
@@
Two typical examples of ansatzes are the UCC ansatz [33, 36, 48, 240] and hardware efficient ansatzes [38]. The UCC is chemically-inspired and thus very expressive for most chemistry and materials, but usually entails a large number of gates and is likely beyond near-term hardware. The hardware-efficient ansatzes are problem-agnostic and designed specifically to use the native gates and connectivity of a given device, but may be impractical to optimize due to the barren plateau problem [241], and also may fail to be sufficiently expressive at low depth, lead to symmetry breaking, and non-smooth potential energy surfaces [242].
@@
Efforts have been put into designing new ansatzes to strike a balance between these two dimensions and unify the advantages of UCC and hardware efficiency [243, 244, 245, 246]. The qubit coupled-cluster ansatz [243, 247, 248] and ADAPT-VQE [246] are among such examples, where a pre-screening process is proposed to select certain Pauli operators to generate the gates in the ansatz, thereby reducing the circuit depth and number of parameters. These methods that trim the number of variational parameters are promising, but there is still a lack of evidence for the scalability of parameter training. An exception may be HMP2 [50], where second-order perturbation theory is employed to reduce the number of variational parameters, while providing real-time, quantum computation-based methodologies to identify additional Pauli operators to include in the VQE progression. These advancements indeed merit further research despite outstanding challenges due to the simplicity of variational quantum algorithm (VQA) in all other respects (particularly in circuit complexity) and the extensive success of classical variational algorithms.
@@
Another challenge with ADAPT-VQE-like algorithms lies in the larger measurement overhead due to the selection of Pauli operators, which can however be mitigated by the use of IC-POVM measurements [249]. While it is important to design an ansatz structure that is most appropriate for a target problem, appropriately initializing the ansatz parameters is an equally critical task when it comes to the trainability of the resulting optimization. This is because random parameter initialization can be extremely far from the optimal parameters, and the ability for VQE to converge to the optimal parameters (and thereby find the ground state) will heavily depend on the optimizer and the effects of noise. Identifying good classical initialization can therefore considerably improve the accuracy and speed of VQE convergence. One promising method is CAFQA [250]. The CAFQA ansatz is built with only Clifford gates and is therefore classically simulable in polynomial time. A discrete optimizer is used to search through different possible Clifford gate parameters until the optimizer converges to a set of parameters that minimizes the VQE expectation (to the best possible extent within the Clifford space). This ‘CAFQA’ state is then used to initialize traditional VQE. There is an opportunity to explore other forms of classical simulation to bootstrap VQE; near-Clifford circuits and tensor network approaches are promising directions [251, 252]. All these methods could benefit from problem-specific optimization. For instance, restricting the CAFQA Clifford search to a state space that is expected to have some overlap with the ground state can improve initialization speed and accuracy.
III.5 Classical data loading
@@
At the highest level, quantum computation involves the following steps. First, a classical description for a specific problem is established. This description is input into a quantum circuit. After executing the quantum circuit, we retrieve classical measurement results as the solution to the problem. Within this workflow, an essential challenge is the data input. A robust quantum algorithm must explicitly detail this input technique and the associated hardware.
@@
For some types of use cases, e.g., quantum simulation algorithms, the problems themselves are quantum mechanical in nature, and often this permits direct mappings to quantum circuits, such as those discussed in Sec. III.1. In more general cases, including when the inputs are classical data, scientists who focus on theoretical complexity analysis often abstract such input methods as an ‘oracle’ and investigate how many times this oracle must be applied, using this as a basis to estimate the algorithm’s complexity. However, creating a physical realization of a quantum oracle is also a fundamental aspect of realizing quantum computation.
@@
QRAM [253] serves as the physical embodiment of the quantum oracle. Unlike traditional RAM, where a single piece of data can be swiftly loaded into the central processor, QRAM allows for multiple classical data items to be loaded concurrently in superposition. This unique capability positions QRAM as a bridge between classical and quantum computing realms. More than that, QRAM can function as a rapid and versatile quantum oracle, effectively addressing the data-input bottleneck prevalent in numerous quantum applications. QRAM is also a fundamental assumption for many quantum algorithms. For instance, many quantum machine learning algorithms [254, 255] may require QRAM as a prerequisite to potentially achieve quantum speedups, although some algorithms have been discovered to achieve the speedups without [256]. Currently, given the significant challenges in physically constructing QRAM, both theoretical and experimental research on QRAM is an active field of study.
@@
For quantum simulation and material sciences, QRAM could primarily be used to provide fast, efficient, and reliable initial states and Hamiltonian encoding. QRAM architectures could provide useful interfaces, benefiting practical use cases of quantum simulation to large-scale, challenging problems in material science.
III.5.1 QRAM construction
@@
  • Quantum Data Center (QDC). Nowadays, data centers are important businesses in the digital society, and it is natural to have a quantum version of data centers in the coming era. Refs. [263, 264] propose a theory of QDCs, where QRAM and quantum networks are combined to provide an elementary definition of QDC. These works show that QDCs with several applications in QRAM as fast quantum memories, will provide fast, secure, and accurate applications in quantum computing, communication, and sensing.
@@
Fault-tolerant designs of quantum random access memory are investigated by [268], showing that QRAM with full fault tolerance might be very challenging on a large scale. On the other hand, noisy QRAM might still be error-resilient for generic setups of decoherence and noisy channels [259]. In fact, though fanout QRAM designs might be vulnerable to decoherence, rendering them non-scalable, the bucket brigade QRAM architecture is known for its high noise resilience [253, 259]. Specifically, the infidelity of a query grows only logarithmically with memory size [259]. Yet, past studies leaned on specific noise models, raising questions about the practical advantages in real-world implementations. Ref. [259] delved into the effects of decoherence on QRAM in a comprehensive manner. A significant takeaway is the affirmation that the logarithmic infidelity scaling persists across varied error channels, including challenges like depolarizing noise and coherent errors. The research identifies limited entanglement among the memory components as the chief reason behind this noise resilience. Interestingly, this insight also suggests potential architectural simplifications without compromising noise resilience, serving an interesting future research question. This understanding implies that with contemporary hardware, QRAM might be potentially realized in real-world noisy environments, facilitating high-fidelity queries even with early fault-tolerant technologies. Additionally, when quantum error correction is integrated, the bucket brigade architecture continues to show promise, enhancing hardware efficiency and fortifying against logical errors [253, 259].
III.5.2 Experimental efforts towards QRAM
@@
III.6 Circuits, transpilation, and architecture
Just as in classical computing, every quantum computation that is to be carried out on a quantum computer needs to be broken down to a netlist of native operations that admit a direct implementation on a quantum computer. One of the most widely-used representation of such a netlist is the so-called quantum circuit, which is to be transpiled for a target quantum computer backend. Different quantum backend hardware may support different qubit-to-qubit connectivity, which limits the interactions between the qubits that may be implemented natively.
To this end, quantum applications in materials science are no exceptions. Indeed, quantum circuits for Heisenberg Hamiltonian simulation [152, 153], relevant for many-body localization and room-temperature superconductivity, and electronic structure calculations for small molecules [48, 50, 53], have been optimized for both pre-fault-tolerant and fault-tolerant settings before. Other examples include particle physics simulations [277, 131, 133], where fault-tolerant cost, including the number of non-Clifford quantum gates and qubit counts, have been minimized. In all these examples, it is critical that the resulting netlist of instructions to be carried out by a quantum computer is optimized, targeting time to solution: a key metric in all types of computation, including quantum. Case in point, if time to solution is of no matter, the computational advantages offered by quantum computing is moot and classical computing suffices, and a high-fidelity operation of a quantum computer would become unnecessary since repeated executions of computation can be performed until an acceptable answer is obtained. The important roles that circuits, transpilation, and architecture will continue to play in leveraging the computational power enabled by quantum computation then form the central topics of this section.
III.6.1 Connectivity and Architecture
We begin by highlighting key topics around qubit connectivity, a fundamental aspect of a quantum processing unit (QPU) that dictates the efficiency, scalability, and feasibility of the quantum devices. We first consider SWAP networks, highlighting their role in enabling Hamiltonian simulation in quantum systems with limited connectivity. The discussion then moves to distributed quantum computing, emphasizing its importance in scaling quantum computing beyond the capabilities of current QPUs. Quantum circuit cutting is also examined for its practicality in decomposing complex quantum circuits into smaller segments, facilitating distributed processing.
@@ III.6.1 Connectivity and Architecture
One of the tools that can be leveraged to effectively implement Hamiltonian simulation primitives on hardware with minimal (linear) connectivity is the SWAP network [278, 279, 280]. Through a SWAP network, depicted in Figure 4, all $\binom{n}{2}\in O(n^{2})$ pairs of (commuting) interactions between $n$ qubits can be accomplished in just $O(n)$ circuit layers, even on a quantum computer architecture that admits nearest-neighbor qubit-to-qubit connectivity. With the exception of an all-to-all architecture that admits a single-instruction implementation of commuting, overlapping two-qubit gates [281, 282, 283], the asymptotic $O(n)$ scaling matches that of an architecture that allows for random-access two-qubit interactions that may be implemented in parallel so long as the interactions do not overlap. In the context of the Materials Science applications in this paper, any “dense” interactions with commuting terms can have a natural implementation on an all-to-all architecture (c.f. Refs. [153, 50, 133, 53]) or on a limited-connectivity architecture with SWAP networks. Reducing the overhead in moving the quantum information around for optimal implementation of a given set of two-qubit interactions remains an area of ongoing research.
@@
Just like classical HPC, distributed quantum computing is necessary for the continued scaling of quantum computing. While monolithic QPUs continue to make significant progress in both sizes and qubit qualities, there remains a huge gap between available QPUs and practical benchmarks. To make matters worse, developments from classical computing demonstrate that the computing and memory capacity gaps between large-scale industrial use cases and single computing cores only get larger, despite having more powerful monolithic cores. To combat this increasing gap, almost all industry-scale classical applications today such as Large Language Model (LLM) training, and even many personal-scale workloads such as intensive gaming, happen in parallel. In fact, distributed computing is only more imperative for quantum computing as single-core scalability poses significant engineering challenges.
There have been distributed quantum computing proposals from both hardware and software researchers. IBM, Quantinuum, and IonQ all announced plans to develop modular QPUs with interconnects to enable distributed quantum computing. Ref. [284] models hardware performances of potential superconducting modular QPUs. Ref. [285] proposes compilation algorithms for distributed QPUs with EPR pair connections. Ref. [286] proposes ClusterVQE which enables domain-specific optimizations to map applications from a target domain (here, VQA) to distributed QPUs. Further research across the stack from applications to devices is necessary to enable distributed quantum computing with tolerable overheads.
@@
Despite its ability to execute previously intractable workloads, there is still a gap between the computational capacity of circuit cutting now and future practical quantum workloads. The first theory proof [287] proposed an exponential classical reconstruction cost. Instead, subsequent works [290] drastically reduce the overhead to sub-exponential but remain prohibitive in the worst case. Filling the gap calls for further developments in classical reconstruction techniques, algorithm-backend co-design, and quantum computing cloud backend system adaptations [291].
@@
The QRAM and QROM memory primitives, discussed earlier in Sec. III.5, can be optimally implemented with appropriate compiler techniques. As one example, consider the Gray Code, a sequence of the $2^{n}$ n-bit bitstrings such that only 1 bit changes between sequential neighbors, for instance, 000, 001, 011, 010, 110, 111, 101, 100. Integrating awareness of the Gray Code in a compiler can significantly reduce gate counts for memory implementations. Additional compilation techniques for these memory primitives, including routing and mapping, are discussed in prior work [292, 274, 275].